dddddddd12:

Which statement about the polynomial x^8 – 26x^4 + 169 is TRUE? A) It has no equal factors. B) It has two equal factors. C) It has four equal factors. D) It has two distinct factors.

dude:

What are the statements?

MoonShadow:

weres the statement

ratnakermehta20:

\[x^{8}-26x^4+169=(x^4)^2-2(x^4)(13)+(13)^2=(x^2-13)^2.\] We conclude that the statement "(B) : It has 2 equal factors." is the right choice.